# Optimal Scaling Quantum Interior Point Method for Linear Optimization

### v1

@@ II Almost-exact Quantum interior point method
where $a_{ij}$ represents the ijij-element of matrix AA. The optimal partition is also defined as
@@
Lemma 1 is a fundamental result in the complexity analysis of IPMs. It means that after a sufficient number of iterations of IPMs, a decision variable can be rounded to zero if it is smaller than $2^{-L}$. Then, by a rounding procedure, one can find an exact optimal solution for linear optimization [roos2005interior, wright1997primal]. In the proposed algorithm, all calculations happen on a quantum machine with precision $\epsilon=2^{-tL}$ where $t$ is a small constant, less than 1010. This high level of accuracy justifies describing the algorithm as almost-exact. The only calculation that happens on a classical computer is updating the solution and vector-vector summation. In this paper, we use the dual logarithmic barrier method, which has a simple framework. At each step of the dual log barrier IPM, we need to solve the following Newton system
@@
One can easily verify that $M=\begin{bmatrix}S^{2}&A^{T}\\ A&0\end{bmatrix}$ is a symmetric positive definite matrix, and so the system (2) has a unique solution [roos2005interior]. Given $s$, one can build block-encodings of implementing matrix $M$ and preparing state $\sigma=\begin{bmatrix}0\\ \frac{1}{\mu}(b-AS^{-1}e)\end{bmatrix}$ efficiently, assuming that matrix AA stored in QRAM in advance. The general steps of the proposed almost exact QIPM using a short-step framework are described in Algorithm 1.
@@ II-A Proof of Theorem 2
It is straightforward to see that $(\tilde{x},y^{k},s^{k})$ is in a $2^{(1-t)L}$-neighborhood of the optimal set for the original problem $(A,b,c)$. As the smallest nonzero element of $s^{*}$ and $x^{*}$ is greater than $2^{-L}$, using partitions $B$ and NN of this solution, by solving a constrained least squares problem, an optimal solution for the original problem can be obtained. For the details of the rounding procedures, refer to Chapter 7 of [wright1997primal].
@@ III Quantum Subroutine
Given AA and $S$ stored on QRAM, the following statements are true:
@@ V Total Complexity
As we can see, the worst-case complexity of IR-AE-QIPM has optimal dimension dependence, $n^{2}$. For the worst-case scenario, storing the dense matrix AA requires ${\cal O}(n^{2})$ operations. Thus, the end-to-end worst-case complexity of an LO solver with respect to dimension can not be less than $n^{2}$.

### v2

@@ Abstract
publicationid: pubid: 979-8-3315-5736-2/25/\$31.00 ©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI 10.1109/QCE65121.2025.00044
@@ II Almost-exact Quantum interior point method
where $a_{ij}$ represents the $ij$-element of matrix $A$. The optimal partition is also defined as
@@
Lemma 1 is a fundamental result in the complexity analysis of IPMs. It means that after a sufficient number of iterations of IPMs, a decision variable can be rounded to zero if it is smaller than $2^{-L}$. Then, by a rounding procedure, one can find an exact optimal solution for linear optimization [roos2005interior, wright1997primal]. In the proposed algorithm, all calculations happen on a quantum machine with precision $\epsilon=2^{-tL}$ where $t$ is a small constant, less than $10$. This high level of accuracy justifies describing the algorithm as almost-exact. The only calculation that happens on a classical computer is updating the solution and vector-vector summation. In this paper, we use the dual logarithmic barrier method, which has a simple framework. At each step of the dual log barrier IPM, we need to solve the following Newton system
@@
One can easily verify that $M=\begin{bmatrix}S^{2}&A^{T}\\ A&0\end{bmatrix}$ is a symmetric positive definite matrix, and so the system (2) has a unique solution [roos2005interior]. Given $s$, one can build block-encodings of implementing matrix $M$ and preparing state $\sigma=\begin{bmatrix}0\\ \frac{1}{\mu}(b-AS^{-1}e)\end{bmatrix}$ efficiently, assuming that matrix $A$ stored in QRAM in advance. The general steps of the proposed almost exact QIPM using a short-step framework are described in Algorithm 1.
@@ II-A Proof of Theorem 2
It is straightforward to see that $(\tilde{x},y^{k},s^{k})$ is in a $2^{(1-t)L}$-neighborhood of the optimal set for the original problem $(A,b,c)$. As the smallest nonzero element of $s^{*}$ and $x^{*}$ is greater than $2^{-L}$, using partitions $B$ and $N$ of this solution, by solving a constrained least squares problem, an optimal solution for the original problem can be obtained. For the details of the rounding procedures, refer to Chapter 7 of [wright1997primal].
@@ III Quantum Subroutine
Given $A$ and $S$ stored on QRAM, the following statements are true:
@@ V Total Complexity
As we can see, the worst-case complexity of IR-AE-QIPM has optimal dimension dependence, $n^{2}$. For the worst-case scenario, storing the dense matrix $A$ requires ${\cal O}(n^{2})$ operations. Thus, the end-to-end worst-case complexity of an LO solver with respect to dimension can not be less than $n^{2}$.